# Can Reasoning LLMs Enhance Clinical Document Classification?11footnotemark: 1

### v1

@@ 5.1 Interpretation of Findings
The findings of this study indicate that reasoning based LLMs generally outperform non-reasoning models in ICD-10 classification tasks, achieving higher accuracy and F1 scores on average. This aligns with the initial hypothesis that explicit reasoning capabilities would enhance the ability of LLMs to capture complex clinical relationships within discharge summaries. However, the results also reveal unexpected patterns. While reasoning models demonstrated superior classification performance, they exhibited lower consistency across repeated experiments compared to non-reasoning models. This suggests that while reasoning based models may improve accuracy, they could also introduce greater variability in results, possibly due to sensitivity to input variations or internal model heuristics. Additionally, while some reasoning LLMs excelled in identifying well-defined medical conditions such as sepsis and myocardial infarction, they struggled with more abstract ICD-10 codes related to patient history and external causes, reinforcing concerns about their contextual understanding. These findings highlight both the potential and limitations of reasoning LLMs, underscoring the need for further refinement to enhance their reliability in clinical applications.
@@ 5.2 Clinical Relevance
The optimization of ICD-10 coding workflows and the broader integration of AI in clinical documentation are of significant clinical relevance. The superior accuracy of reasoning based LLMs suggests that they could serve as valuable tools for assisting clinical coders in accurately classifying discharge summaries, potentially reducing manual workload and minimizing coding errors. However, the observed variability in their consistency raises concerns about reliability, highlighting the importance of model calibration and validation before deployment in real world settings. Additionally, the inability of some models to correctly classify abstract or context dependent ICD-10 codes underscores the need for further refinements, such as enhanced domain specific training and better handling of nuanced medical contexts. If properly fine tuned and integrated, these AI driven coding assistants could streamline the ICD-10 classification process, improve coding accuracy, and ultimately support more efficient clinical documentation and billing workflows, leading to better healthcare data quality and decision making.
@@ 6 Conclusion
This study also has significant implications for the future of automated clinical coding and decision support systems. The superior consistency of non-reasoning models makes them well suited for routine, high volume ICD-10 coding tasks, where pattern recognition and statistical learning ensure accuracy and efficiency. Their stable performance across repeated trials supports large scale deployment, reducing manual coding and minimizing human error. Meanwhile, the strong accuracy and F1 scores of reasoning models suggest their value in handling ambiguous cases requiring deeper contextual understanding. A hybrid approach leveraging non-reasoning models for routine coding and reasoning models for complex cases could enhance both efficiency and accuracy. Additionally, the consistent classification of non-reasoning models highlights their potential for real-time decision support, such as identifying diagnostic trends or recommending treatments. Integrating both models into a unified system could improve coding accuracy, streamline workflows, and enhance clinical decision making, ultimately benefiting patient care and reducing administrative burden.

### v2

@@ 5.1 Interpretation of Findings
The findings of this study indicate that reasoning LLMs generally outperform non-reasoning models in ICD-10 classification tasks, achieving higher accuracy and F1 scores on average. This aligns with an initial hypothesis that explicit reasoning capabilities would enhance the ability of LLMs to capture complex clinical relationships within discharge summaries. However, the results also reveal unexpected patterns. While reasoning models demonstrated superior classification performance, they exhibited lower consistency across repeated experiments than non-reasoning models. This suggests that while reasoning models may improve accuracy, they could also introduce greater variability in results, possibly due to sensitivity to input variations or internal model heuristics. Additionally, while some reasoning LLMs excelled in identifying well-defined medical conditions such as sepsis and myocardial infarction, they struggled with more abstract ICD-10 codes related to patient history and external causes, reinforcing concerns about their contextual understanding. These findings highlight both the potential and limitations of reasoning LLMs, underscoring the need for further refinement to enhance their reliability in clinical applications.
@@ 5.2 Clinical Relevance
The optimization of ICD-10 coding workflows and the broader integration of AI in clinical documentation are of significant clinical relevance. The superior accuracy of reasoning LLMs suggests that they could serve as valuable tools for assisting clinical coders in accurately classifying discharge summaries, potentially reducing manual workload and minimizing coding errors. However, the observed variability in their consistency raises concerns about reliability, highlighting the importance of model calibration and validation before deployment in real world settings. Additionally, the inability of some models to correctly classify abstract or context-dependent ICD-10 codes underscores the need for further refinements, such as enhanced domain-specific training and better handling of nuanced medical contexts. If properly fine-tuned and integrated, these AI-driven coding assistants could streamline the ICD-10 classification process, improve coding accuracy, and ultimately support more efficient clinical documentation and billing workflows, leading to better healthcare data quality and decision making.
5.3 Future Work